OpenStudy (anonymous):

Solve for x a) 3^x= 300 b) 30*5^1-x = 0.15 c) 3^x+2 = 2^1-x

OpenStudy (anonymous):

b) 30 * 5^1- x = 0.15 30 * 5 - x= 0.15 150 - x = 0.15 -x = 0.15 - 150 = -149.85 x = 149.85

OpenStudy (anonymous):

Sorry, couldn't get the other answers.

OpenStudy (anonymous):

hmm i think your supposed to use use logarithms??-- but thanx anyways!